OpenStudy (anonymous):

Solve the recurrence relation using iteration: $$a_n = a_{n-1} + 1 + 2^{n-1}\\ a_0 = 0$$ So $a_1$ = 2, $a_2$ = 5... I really have no idea how to solve recurrence relations that aren't of the linear, homogeneous type with constant coefficients. How would I solve this one? by iteration? to *n*? It's undefined... What are the properties of this recurrence relation (linear, homogeneous, etc.) and why? Then I'm supposed to be able to prove the solution to this by mathematical induction, I guess since it's solved "using iteration"... I don't know how to do this either.

OpenStudy (anonymous):

Solve the recurrence relation using iteration: \(a_n = a_{n-1} + 1 + 2^{n-1}\\ a_0 = 0\) So \(a_1\) = 2, \(a_2\) = 5... I really have no idea how to solve recurrence relations that aren't of the linear, homogeneous type with constant coefficients. How would I solve this one? by iteration? to *n*? It's undefined... What are the properties of this recurrence relation (linear, homogeneous, etc.) and why? Then I'm supposed to be able to prove the solution to this by mathematical induction, I guess since it's solved "using iteration"... I don't know how to do this either.

OpenStudy (anonymous):

i cheated and found the answer via wolfram wondering what would happen if the \(2^{n-1}\) was not there i.e. if it was just \[a_n=a_{n-1}+1\] would the answer be \[a_n=c+n\]?

OpenStudy (anonymous):

oh yes, clearly it would be that

OpenStudy (anonymous):

now with the \(2^{n-1}\) there, i think you just kick the power up by one and get \[2^n\]

OpenStudy (anonymous):

or maybe \[2^n-1\] since you are summing \[\sum 2^{n-1}\]

OpenStudy (anonymous):

try \[a_n=n+2^n\]

OpenStudy (anonymous):

\[a_0=0\] so i guess that is wrong, try \[a_n=n+2^n-1\]

OpenStudy (anonymous):

Oh that looks good!! thank you!! but how are we supposed to get this, just by guessing? I was trying out similar formulas but hadn't quite gotten that one yet.